hi i think im close but still cant get this one.
sum of three numbers =129
second number is 9 less the first
third number is 3 times the second
1st #=
2nd #=
3rd #=

Answer by Boreal(15235)

x, y, and z are the numbers
x+y+z=129
y=x-9
z=3(x-9)
substitute
x+x-9+3x-27=129
5x-36=129
5x=165
x=33
the numbers are 33, 24, and 72, and they add to 129.